Ball landed on the ground with impact velocity of 30 m/s.

a.) from how high it was dropped?
b.) where was it at time= 0.5s?
c.) how long did it take to fall?
1
Expert's answer
2019-10-23T09:53:36-0400

The final velocity of the ball is "v_f=30\:\rm m/s," the acceleration due gravity "g=9.8\:\rm m/s^2."

(a) Height


"H=\frac{v_f^2}{2g}=\frac{30^2}{2\times 9.8}=46\:\rm m"

(b) Position at 0.5s


"y=H-\frac{gt^2}{2}=46-\frac{9.8\times 0.5^2}{2}=44.8\:\rm m\quad (above \; the \; ground)"

(c) Time of motion


"t=\frac{v_f}{g}=\frac{30}{9.8}=3\:\rm s"